answers asappp

selected for Flow velocity 4 m/s 4. Give sound velocity for carbon dioxide at temperature 800 C degree.
Answer #1

yu speed of sound - SORT N=13 (coz is triatsmic) R=&= 8.31% = 0.434 kJ / 1g k=1345 /gk. T= 800 c. 10738 speed to sound 778 mi